OpenStudy (anonymous):

For the function f(x) = 6x − 8, identify the values of f(0), f(1/2), and f(−2).

OpenStudy (theeric):

Hello! Do you know how functions work, @hyesukm52 ?

OpenStudy (anonymous):

No... :(

OpenStudy (theeric):

\(f\) is the function's name, and the parentheses hold the independent variable. Independent variables are just themselves - they don't \(\text{depend}\) on anything else. So, this independent variable just puts itself in the equation. The function works but putting numbers in the function's independent variable's place. For example, \(f(5)=6(5)-8\\f(4)=6(4)-8\\f(300)=6(300)-8\) And you can compute these: \(f(5)=6(5)-8=22\\f(4)=6(4)-8=16\\f(300)=6(300)-8=1792\)

OpenStudy (theeric):

So try to compute what is indicated by your values! Good luck!